OpenStudy (mackenzie2013):

A rectangular prism is 3 meters long, 4 meters wide, and has a total surface area of 94 square meters. What is its height? Help ! Please Explain Steps !

OpenStudy (ddcamp):

SA = 2lw + 2wh + 2lh SA = 2(3)(4) + 2(4)h + 2(3)h SA = 24 + 14h Think you can solve it from here?

OpenStudy (anonymous):

plug in 94 in place of SA and then solve for h:\[94=24+14h\]

OpenStudy (mackenzie2013):

Its multiple choice, Im not getting any of the answers . 3 m 4 m 5 m 6 m

OpenStudy (anonymous):

that's great, now solve to find the answer